Explain why lines m and n are parallel. V-axis JAM - 1,5) (9, 4) (60) X-axis VV
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case the answer has helped you


Slope of m is (9-5)/(4-1) = 4/3

Slope of n is (4-0)/(9-6) = 4/3

So, the slopes are equal, meaning m and n are parallel